Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
bersanetti, magazzu, nardecchia - 16:27 Wednesday 29 October 2025 (68058) Print this report
Comment to NI absorption measurement (68051)

To investigate the hypothesis that the structure in the upper part of the HWS map is caused by the green beam, Diego modified the automation configuration to keep it OFF after the unlock. We then performed a new measurement with HWS-DET. For this purpose, I switched OFF the HWS-INJ SLED and switched ON the HWS-DET SLED.

The ITF was locked at CARM NULL 1F at 11:04 UTC. Acquisition started at 14:28 UTC (folder 20251029T1528), and the ITF gently unlocked at 14:36 UTC (WF14).

The structure was also present in the maps (see figure).
The measurement was interrupted, and the lock acquisition was restarted.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×